OpenStudy (anonymous):

A theme park charges $52 for a day pass and $110 for a week pass. Last month, 4,432 day passes were sold and 1,979 week passes were sold. Which is the closes estimate of the total amount of money paid for the day and week passes for last month? A. $200,000 B. $300,000 C. $400,000 D. $500,000

OpenStudy (anonymous):

Round the numbers: 50 x 4500 + 100 x 2000 = 225000 + 200000 = 425000 = 400, 000 dollars